OpenStudy (anonymous):

prove sinh x + sinh y = 2*sinh((x + y)/2)*cosh((x - y)/2) this is what i done so far Let a=(x+y)/2 = 2a b=(x-y)/2 =2b 2a+2b=x 2a-2b=y

OpenStudy (irishboy123):

if you do the RHS as \[\large 2 \left( \dfrac{e^{\frac{x+y}{2}} - e^{-\frac{x+y}{2}}}{2}\right).\left( \dfrac{e^{\frac{x-y}{2}} + e^{-\frac{x-y}{2}}}{2}\right)\]

OpenStudy (anonymous):

did that but i stuck halfway, that's why i try to apply the identities

OpenStudy (er.mohd.amir):

multiply that term by term

OpenStudy (irishboy123):

well, first term of the expansion would be \[\large e^{\frac{x+y}{2}}.e^{\frac{x-y}{2}} = e^x\] it just unfolds

OpenStudy (anonymous):

so i got \[2(\frac{ e^x + e^y- e^y-x}{2 })\]

OpenStudy (er.mohd.amir):

you miss one 2 in denominater

OpenStudy (anonymous):

oh yeah i forgot to multiply it

OpenStudy (anonymous):

after this i should apply the exponent properties, but how can i can make x and y become negative

OpenStudy (er.mohd.amir):

e^x + e^y -(e^-x) - (e^-y) upon 2

OpenStudy (anonymous):

i see my mistake now, thank you for your helps, i really appreciate it